# 208Pb nuclear charge radius revisited: closing the fine-structure-anomaly gap

### v1

@@ Abstract
A comprehensive reevaluation of the root-mean-square nuclear charge radius is presented for the doubly magic 208Pb extracted from muonic spectroscopy measurements. By integrating rigorous theoretical quantum electrodynamics calculations, state-of-the-art numerical methods, and a systematic reanalysis of the uncertainties, we reduced the long-standing muonic fine-structure anomaly and improved the goodness of fit by a factor of twenty. The resulting value of 5.5062(5) fm is fairly consistent with the previously reported muonic spectroscopy value, and three standard deviations larger than the commonly used compilation data, which indicates that the current value and its uncertainty could be significantly underestimated. Our study paves a new path for systematic reevaluation of all rms radii based on muonic spectroscopy.
@@
Accurate calculations of nuclear parameters from first principles are particularly challenging for heavy nuclei. Since direct measurements of nuclear charge radii are not possible, indirect extraction with significant theory input is necessary. Muonic atoms are highly sensitive to these nuclear parameters, such that the most accurate method to measure rms charge radii is based on muonic atom spectroscopy [1, 2]. It is also common to combine the values obtained from muonic atoms with electron scattering cross-section data [3] via Barrett moments [4].
Motivated by the success of light muonic atom spectroscopy [5, 6, 7] and new measurements of heavy muonic atoms with microtargets [8], we started theoretical efforts to extract the rms radii of heavy atoms. A literature study revealed the existence of another puzzle, the muonic fine-structure anomaly, discovered about 50 years ago [9, 10, 11, 12] and not mentioned since then: the theoretically predicted spectra of lead, tin, and zirconium exhibit small but systematic deviations from the experimentally measured values, for any value of the assumed nuclear rms charge radius. This discrepancy was attributed to the least understood and most complicated of all contributions: nuclear polarization. However, our rigorous microscopic calculations of the nuclear polarization have shown that it is most likely not solely responsible for the anomaly [13]. This aligns with the poor fit quality addressed above and hints at the uncertainty of the nuclear radius of double-magic 208Pb being significantly underestimated. The same situation applies to other muonic atoms, as the methods for rms charge radius extraction have been the same.
@@
Theory improvements. — We calculated all individual contributions to the muonic spectrum which contribute $\gtrsim$10of the experimental precision. The leading contribution is the finite nuclear size (FNS) effect, obtained by numerically solving the Dirac equation with a nuclear potential derived from a two-parameter Fermi (2pF) nuclear charge model (see Sup. Mat. for details).
State FNS Ue WK KS $\mu$Ue HadUe SE NP RRecoil Screening
@@
$3d_{5/2}$ 12121.477 19.848 -0.185 0.071 0.001 $0$0.070(5) 0.037(37) -1.143(0) 5.523
@@
In addition to the electron-positron loop, VP can also originate from any other particle-antiparticle (e.g. muon-antimuon) loop. However, these effects are exponentially suppressed by the mass of the virtual particles; thus, going beyond leading order is only necessary for electronic Ue at the current level of accuracy. The muon-antimuon loop gives the muonic Ue ($\mu$Ue) correction, which can be calculated by a straightforward replacement of the virtual particle mass. For the hadronic Ue (HadUe) contribution, we employed the approach [22] derived from hadronic scattering data, and adapted it for muonic atoms [23].
@@
Experimental data. — For our current evaluation, we have used the experimental measurements of $\mu-^{208}$b reported in Ref. [11] (see Table 2) and recently confirmed them to be reliable [28]. The data are split in three runs, which correspond to the high-, medium-, and low-energy parts of the spectrum. This split was arranged to allow for specific tuning of the detector. We performed the fit with respect to all transitions, but also analyzed the results from individual runs.
@@
Results and Discussion. — The best-fit 2pF parameters $c$ and $a$, together with the corresponding rms charge radius, are shown in Table 3. Fig. 2 depicts the likelihood contours for the 2pF parameters up to the $5\sigma$ level. The fit was performed for each of the 3 runs separately and the final value was extracted from a combined fit to all the experimental data. The individual fits are represented by different colours: purple, green, and blue correspond to runs I, II, and III, respectively. The full fit is depicted in orange. The insets show the residual transition-energy differences between best-fit theory and measurement for each of the fits. The ellipses corresponding to the 3 individual runs are in mild tension with each other which is ultimately responsible for the poor fit quality of the full fit ($\chi^{2}/\mathrm{DoF}=9.7$). Interestingly, the best-fit values for the individual runs still fall closely to a line of constant rms charge radius. Note that the figure does not show the best-fit value for Run I as it is far off the other two, but can be found in Table 3.
@@
Figure 2: The fit of the spectrum of $\mu-^{208}$b to a 2pF distribution. The colored regions show the $5\sigma$ likelihood for the 2pF parameters and the dotted contours highlight the $5$, $3$, and $1\sigma$ steps. Each colour corresponds to a fit of different groups of transitions: purple only fits Run I, green corresponds to Run II, blue to Run III, and orange is the combined fit of all the transitions. The insets show the resulting residual transition energies in the same colours, with measured transitions in darker shade and resulting fine-structure splittings in lighter colour. The previous fit as taken from Ref. [11] is shown in black wherever available, and the error bars include experimental errors only as the theory uncertainty was not explicitly quoted. The dashed lines correspond to curves of constant radius. The coloured circles indicate the best-fit parameters and the red circle the final values for the overall fit. Note that the best-fit value for Run I lies outside of the plotted parameter range, as indicated by the red arrow.
@@
Contrary to conventional understanding, our analysis demonstrates that transitions involving $2s$, $2p$ and $3p$ states impose at least equally stringent constraints on radius extraction as traditionally emphasized $2p-1s$ transitions, which can be important for future experiments.
Our final value is consistent with the previous work in Ref. [11], which analyzed the same muonic spectral data, but lacks rigorous QED predictions for some of the effects. Compared to the accepted value, which combines muonic spectroscopy with electron-scattering measurements, we find a significantly larger rms charge radius. We are convinced that the current standard value might be too small and requires reconsideration in light of this work. Our results strongly indicate the necessity of a systematic reevaluation of the rms charge radii across the periodic table.
Acknowledgements. — The Authors thank A. Knecht, B. Ohayon, V. A. Yerokhin, and V. A. Zaytsev for discussions. N. S. O thanks ISOQUANT for financial support. This article comprises parts of the PhD thesis work of Z. M. and Z. S. to be submitted to Heidelberg University.

### v2

@@ Abstract
A comprehensive reevaluation of the root-mean-square nuclear charge radius is presented for the doubly magic 208Pb extracted from muonic spectroscopy measurements. By integrating rigorous theoretical quantum electrodynamics calculations, state-of-the-art numerical methods, and a systematic reanalysis of the uncertainties, we reduced the long-standing muonic fine-structure anomaly and improved the goodness of fit by a factor of twenty. The resulting value of 5.5062(5) fm for a Fermi distribution is fairly consistent with the previously reported muonic spectroscopy value, and three standard deviations larger than the commonly used compilation data, which indicates that the current value and its uncertainty could be significantly underestimated. Attributing the remaining discrepancy to theory errors which can not be rigorously calculated we suggest the rms charge radius with reduced model dependence to be 5.5062(17) fm. This work sets an improved benchmark for charge radius extraction in heavy nuclei and paves a path for systematic reevaluations across the nuclear chart.
@@
Accurate calculations of nuclear parameters from first principles are particularly challenging for heavy nuclei. Since direct measurements of nuclear charge radii are not possible, indirect extraction with significant theory input is necessary. Muonic atoms are highly sensitive to these nuclear parameters, such that the most accurate method to measure rms charge radii is based on muonic atom spectroscopy [1, 2]. Motivated by the success of light muonic atom spectroscopy [3, 4, 5] and new measurements of heavy muonic atoms with microtargets [6], we started the development of a significantly more accurate theory for the rms radii extraction of heavy nuclei.
While muonic atom spectroscopy provides the most precise determinations of rms charge radii, its application to heavy nuclei has revealed a persistent discrepancy, namely the muonic fine-structure anomaly [7, 8, 9, 10], first observed nearly 50 years ago but never resolved. The theoretically predicted spectra of lead, tin, and zirconium exhibit small but systematic deviations from the experimentally measured values, for any value of the assumed nuclear rms charge radius. This discrepancy was attributed to the least understood and most complicated of all contributions: nuclear polarization. However, our rigorous microscopic calculations of the nuclear polarization have shown that it is most likely not solely responsible for the anomaly [11]. This aligns with the poor fit quality addressed above and hints at the uncertainty of the nuclear radius of double-magic 208Pb being significantly underestimated. The same situation applies to other muonic atoms, as the methods for rms charge radius extraction have been the same.
@@
Theory improvements. — We calculated all individual contributions to the muonic spectrum which contribute $\gtrsim$10% of the experimental precision. The leading contribution is the finite nuclear size (FNS) effect, obtained by numerically solving the Dirac equation with a nuclear potential derived from a two-parameter Fermi (2pF) nuclear charge model (see Sup. Mat. for details).
[…]
@@
$3d_{5/2}$ 12121.477 19.848 -0.185 0.071 0.001 0.000 -0.070(5) 0.037(37) -1.143(0) 5.523
Table 1: Corrections to muonic $^{208}\mathrm{Pb}$ binding energies (keV) calculated perturbatively with FNS wavefunctions from the 2pF model for the best-fit values $a=0.5283$ fm, $c=6.6409$ fm, $\text{rms radius}=5.5062$ fm. A positive correction refers to an increase in binding energy.
@@
Experimental data. — For our current evaluation, we have used the experimental measurements of $\mu-^{208}$Pb reported in Ref. [9] (see Table 2) and recently confirmed them to be reliable [27]. The data are split in three runs, which correspond to the high-, medium-, and low-energy parts of the spectrum. This split was arranged to allow for specific tuning of the detector. We performed the fit with respect to all transitions, but also analyzed the results from individual runs.
@@
Results and Discussion. — The best-fit 2pF parameters $c$ and $a$, together with the corresponding rms charge radius, are shown in Table 3. Fig. 2 depicts the likelihood contours for the 2pF parameters up to the $5\sigma$ level. The fit was performed for each of the 3 runs separately and the final value was extracted from a combined fit to all the experimental data. The individual fits are represented by different colours: purple, green, and blue correspond to runs I, II, and III, respectively. The full fit is depicted in orange. The insets show the residual transition-energy differences between best-fit theory and measurement for each of the fits. The ellipses corresponding to the 3 individual runs are in mild tension with each other which is ultimately responsible for the poor fit quality of the full fit ($\chi^{2}/\mathrm{DoF}=9.5$). Interestingly, the best-fit values for the individual runs still fall closely to a line of constant rms charge radius. Note that the figure does not show the best-fit value for Run I as it is far off the other two, but can be found in Table 3.
@@
Even though the theory uncertainties have been estimated to the best of our ability, the quality of those estimates can not be checked directly. Assuming that the experimental data is correct and the spectrum is fully described within QED we attribute the remaining discrepancy to unaccounted uncertainties like, e.g. model dependence. By normalizing the error such that the $\chi^{2}/\rm{DoF}$ becomes unity, and ensuring that the spectrum is reproduced within the accuracy of the calculation, we arrive at a conservative, reduced model-dependent rms charge radius of $5.5062(17)$ fm.
Figure 2: The fit of the spectrum of $\mu-^{208}$Pb to a 2pF distribution. The colored regions show the $5\sigma$ likelihood for the 2pF parameters and the dotted contours highlight the 55, 33, and $1\sigma$ steps. Each colour corresponds to a fit of different groups of transitions: purple only fits Run I, green corresponds to Run II, blue to Run III, and orange is the combined fit of all the transitions. The insets show the resulting residual transition energies in the same colours, with measured transitions in darker shade and resulting fine-structure splittings in lighter colour. The previous fit as taken from Ref. [9] is shown in black wherever available, and the error bars include experimental errors only as the theory uncertainty was not explicitly quoted. The dashed lines correspond to curves of constant radius. The coloured circles indicate the best-fit parameters and the red circle the final values for the overall fit. Note that the best-fit value for Run I lies outside of the plotted parameter range, as indicated by the red arrow.
@@
Conclusions. — We performed fully rigorous, state-of-the-art QED calculations for the spectrum of muonic lead to extract the nuclear rms charge radius from a fit to experimental measurements. Using a two-parameter Fermi distribution, we determine the nuclear rms radius value of 5.5062(5) fm, achieving a twentyfold improvement in fit quality over previous analyses. The remaining overall high $\chi^{2}/{\rm DoF}=9.5$ we traced back to an inconsistency between the fits of groupings of the transitions by energy.
@@
To reduce residual model dependence, we conservatively normalize uncertainties to achieve $\chi^{2}/\mathrm{DoF}=1$, yielding a reduced model-dependent estimate of 5.5062(17) fm. This value is consistent with the previous work in Ref. [9], which analyzed the same muonic spectral data, but our work incorporates complete QED corrections for the first time. Our result is significantly larger than the currently accepted standard based on combined muonic and electron-scattering data.
We are convinced that the current standard rms charge radius of 208Pb may be underestimated and requires reconsideration in light of this work. More broadly, our results strongly indicate the necessity of a systematic, theory-driven reevaluation of the rms charge radii across the periodic table along the avenue demonstrated here using muonic spectroscopy and modern QED theory.
Acknowledgements. — The Authors thank A. Knecht, F. Heiße, B. Ohayon, V. A. Yerokhin, and V. A. Zaytsev for discussions. This work is part of and funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the Collaborative Research Centre, Project-ID No. 273811115, SFB 1225 ISOQUANT. This article comprises parts of the PhD thesis work of Z. M. and Z. S. to be submitted to Heidelberg University.